# RoboReflect: Robotic Reflective Reasoning for Grasping Ambiguous-Condition Objects

### v1

@@ IV-B Overall Performance
To demonstrate the advantages of our framework, we compare it with AnyGrasp [6] as a baseline and a high-level policy planner using GPT-4V [7], which excludes other framework components. We evaluate the performance of RoboReflect against baseline methods and compare their grasping success rates on the eight objects, as shown in Table I. The results indicate that both AnyGrasp and GPT-4V performed poorly when dealing with ambiguous condition objects, with low success rates, demonstrating that this category of objects presents a challenge for general grasping algorithms. RoboReflect shows a significant advantage compared to the baseline methods. Compared to the second-best method, RoboReflect improves the average accuracy by 21.25%; compared to GPT-4V, it improves by 50%. These results highlight the exceptional performance of RoboReflect when grasping ambiguous conditions objects.
@@
When grasping a partially empty tissue bag, the top half is prone to deformation, causing grasping failures regardless of how AnyGrasp or GPT-4V attempts to handle it. However, the extremely high success rate (70%) of RoboReflect highlights the superiority of our framework. When grasping cookies, external hard drives and ice cream bars present greater challenges to the baseline methods. The grasping of these objects contains some ambiguous conditions, which may be common sense to humans, but difficult to understand for baseline methods. Therefore, the success rate of AnyGrasp and GPT-4V is extremely low, as 10%, 0%, 0% and 20%, 30%, 30%, respectively. In contrast, RoboReflect can overcome these ambiguous conditions through autonomously reflective, thus achieving correct grasping with high performance, 60%, on three objects.
It’s worth noting that when grasping cup noodles and closed-lid cups, the performance of RoboReflect is not much different from that of AnyGrasp and GPT-4V, with similar success rates. This is because these two objects are in an ideal state and not prone to characteristic changes, so there is no need to think too much about the grasping approach. However, when grasping open-lid cups and open-lid cup noodles, the success rate of AnyGrasp has a similar performance to RoboReflect, but much higher than GPT-4V. This is because AnyGrasp initially tends to grasp cup-like objects from the body of the cup, while our method initially grasps from directly above and needs to reflect and reason to master the correct grasping method. We also show some qualitative results of RoboReflect in Fig. 4.
@@ V CONCLUSION
In this paper, we introduced RoboReflect, a novel framework for grasping robots that enables autonomous reflection and action correction using large vision-language models (LVLMs). By addressing the challenges posed by ambiguous-condition objects in complex environments, RoboReflect enhances robotic grasping performance through continuous reflection and strategy refinement. Our experiments on eight common objects across three categories demonstrate that RoboReflect significantly improves grasp success rates compared to existing methods, such as AnyGrasp and GPT-4V. The framework’s capability to store successful strategies for future reference further strengthens its capability to handle diverse scenarios without human intervention. These results highlight the importance of incorporating autonomous reflection into robotic systems, offering a promising solution for more resilient and adaptable robots in real-world applications.

### v2

@@ IV-B Overall Performance
To demonstrate the advantages of our framework, we compare it with ReKep [6], AnyGrasp [7] as baselines and a high-level policy planner using GPT-4V [8], which excludes other framework components. We evaluate the performance of RoboReflect against baseline methods and compare their grasping success rates on the eight objects, as shown in Table I. The results indicate that both ReKep, AnyGrasp and GPT-4V performed poorly when dealing with ambiguous condition objects, with low success rates, demonstrating that this category of objects presents a challenge for general grasping algorithms. RoboReflect shows a significant advantage compared to the baseline methods. Compared to the AnyGrasp, RoboReflect improves the average accuracy by 21.25%; compared to GPT-4V, it improves by 50% and compared to ReKep it improves by 17.5%. These results highlight the exceptional performance of RoboReflect when grasping ambiguous conditions objects.
@@
When grasping a partially empty tissue bag, the top half is prone to deformation, causing grasping failures regardless of how AnyGrasp or GPT-4V attempts to handle it. However, the extremely high success rate (70%) of RoboReflect highlights the superiority of our framework. When grasping cookies, external hard drives and ice cream bars present greater challenges to the baseline methods. The grasping of these objects contains some ambiguous conditions, which may be common sense to humans, but difficult to understand for baseline methods. Although ReKep demonstrates the capability to handle basic object manipulation tasks, such as grasping ice cream, its accuracy of 50% remains slightly inferior to that of our proposed method, its grasping performance is significantly limited when faced with objects that require more understanding, such as cookies and external hard drives, with a success rate of only 10%. Therefore, the success rate of AnyGrasp and GPT-4V is extremely low, as 10%, 0%, 0% and 20%, 30%, 30%, respectively. In contrast, RoboReflect can overcome these ambiguous conditions through autonomously reflective, thus achieving correct grasping with high performance, 60%, on three objects.
It’s worth noting that when grasping cup noodles and closed-lid cups, the performance of RoboReflect is not much different from that of ReKep, AnyGrasp and GPT-4V, with similar success rates. This is because these two objects are in an ideal state and not prone to characteristic changes, so there is no need to think too much about the grasping approach. However, when grasping open-lid cups and open-lid cup noodles, the success rate of ReKep and AnyGrasp has a similar performance to RoboReflect, but much higher than GPT-4V. This is because AnyGrasp initially tends to grasp cup-like objects from the body of the cup, while our method initially grasps from directly above and needs to reflect and reason to master the correct grasping method. To present the experimental results more intuitively, Fig. 3 shows the comparison of the grasping pose generation effects of AnyGrasp, ReKep and our method on these objects. We also show some qualitative results of RoboReflect in Fig. 4.
@@ V CONCLUSION
In this paper, we introduced RoboReflect, a novel framework for grasping robots that enables autonomous reflection and action correction using large vision-language models (LVLMs). By addressing the challenges posed by ambiguous-condition objects, RoboReflect enhances robotic grasping performance through continuous reflection and strategy refinement. Our experiments on eight common objects across three categories demonstrate that RoboReflect significantly improves grasp success rates compared to existing methods. The framework’s capability to store successful strategies for future reference further strengthens its capability to handle diverse scenarios without human intervention. These results highlight the importance of incorporating autonomous reflection into robotic systems, offering a promising solution for more resilient and adaptable robots in real-world applications.